# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

### v5

@@ 6. Discussion
We discuss the implications of our estimates. In particular, we assess the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data. Our estimates suggest that while both data sets are spatially representative, they miss the majority of street vendors. We provide several explanations for this discrepancy.
@@
We believe this application highlights the value of modeling respondents as a point process. We believe our approach is relevant for a wide variety of surveys in which units cannot be chosen at random from a sampling frame. For example, individuals may be sampled on a beach or at various entrances of a park. Future work might explore additional point process models that describe the behavior of the population. For example, we have found that when vendors cluster within markets, standard errors become overdispersed. However, it is possible that vendors protect their turf, leading to underdispersion. It may even be that some vendors benefit from operating near each other, while others benefit from maintaining their distance.

### v6

@@ 6. Discussion
We conclude with a discussion. In Section 6.1, we consider how the theoretical justification provided in Section 3.2 and extended in Sections 4.1 and 4.2 might be further extended to accommodate more complex models. In Section 6.2, we consider some implications of the estimates provided in Section 5.
6.1 Theoretical Extensions
In Section 3.2, we develop a theoretical justification for ratio estimation based on the theory of point processes. We assume the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes. Our approach yields simple, closed-form expressions for the estimates and standard errors on which policy can be based.
In Section 4, we relax both assumptions to determine the sensitivity of these expressions. Our relaxations yield additional closed-form expressions for the estimates and standard errors. These expressions are more complicated than those derived in Section 3.2. While the estimates and standard errors are not identified from the survey data alone, they provide insight into how exactly a violation would change an estimate or standard error, again facilitating policy.
The results of Section 5 suggest that deviations from our assumptions do not materially change our findings, except in extreme cases that we do not believe hold in practice such as if the response probability were inversely proportional to enforcement. However, it is possible our findings would change if we were to relax both assumptions further. For example, in Section 4.1 we relax the assumption that the distribution of respondents is representative by generalizing the Poisson process to accommodate probability weights with a known mean and covariance. Future work might consider the more general case in which the mean and covariance of the weights are estimated. One might model the response probability with additional data using logistic regression from which the mean and covariance of the weights could be estimated. The additional uncertainty would inflate the standard errors relative to Section 4.1. See Brick and Montaquila (2009, sec. 3.2) for additional considerations when modeling the response probability.
For another example, in Section 4.2 we relax the assumption that the spatial distribution is well-approximated by a set of inhomogeneous Poisson processes. We consider the possibility that the spatial distribution of vendors is better approximated by a set of inhomogeneous negative binomial process, which produces overdispersion when many vendors cluster within a market. Future work might consider even more general point process models. For example, one might model the location of vendors as a Neyman-Scott or Cox process, which also produce overdispersion. Overdispersion is not inevitable, however, as it is possible that vendors protect their turf, leading to underdispersion. It may even be that some vendors benefit from operating near each other, while others benefit from maintaining their distance. See Cressie and Wikle (2011, chap. 4.3) for additional considerations when modeling Cox or dependent processes.
Finally, we assume that spatial regions are sufficiently large or dense that inference can be achieved using asymptotic arguments. Information is borrowed across subregions through estimation of a shared ratio. Future work might compare this approach with other methods for sharing information between subregions. For example, the literature on small area estimation often uses hierarchical models such as the one specified in Section 7.6 to borrow information across subregions, often called subdomains. See Rao (2003) for for additional considerations when estimating counts in small regions. One way to augment the hierarchical approach outlined in Section 5.2 is to model the $\mu_{i}$ as a function of auxiliary data, such as those described in that section.
6.2 Practical Implications
We consider the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data. Our estimates suggest that while both data sets are spatially representative, they miss the majority of street vendors. We provide several explanations for this discrepancy.
@@ 6.2 Practical Implications
Even if the vendor responds to the ACS, they may fail to indicate that they are a street vendor or work in New York City. Indeed, approximately 25% of individuals classified as Door-To-Door Sales Workers, News And Street Vendors, And Related Workers with place of work in New York City had their occupation imputed due to missing values as indicated by the occupation allocation flag. Even more, 40%, had their work location imputed as indicated by the place of work allocation flag. These imputation rates are high but not uncommon. Approximately a third of individuals classified as working in New York City had an occupation with a higher allocation rate, while 10% had a higher place of work allocation rate.